a line passes through (3a,4) and (-1,2) and is parallel to -4x +2y= 6 find a
Answer by edjones(8007)

(3a,4) and (-1,2)
-4x +2y= 6
.
2y=4x+6
y=2x+3 the slope, m=2.
.
(4-2)/(3a-(-1))=2
2/(3a+1)=2
2/1=2 so 3a+1=1 and a=0
.
Ed